# Detecting Cognitive Impairment and Psychological Well-being among Older Adults Using Facial, Acoustic, Linguistic, and Cardiovascular Patterns Derived from Remote Conversations

### v1

@@ Abstract
Abstract
INTRODUCTION: The aging society urgently requires scalable methods to monitor cognitive decline and identify social and psychological factors indicative of dementia risk in older adults.
METHODS: Our machine learning models captured facial, acoustic, linguistic, and cardiovascular features from 39 individuals with normal cognition or Mild Cognitive Impairment derived from remote video conversations and classified cognitive status, social isolation, neuroticism, and psychological well-being.
RESULTS: Our model could distinguish Clinical Dementia Rating Scale of 0.5 (vs. 0) with 0.78 area under the receiver operating characteristic curve (AUC), social isolation with 0.75 AUC, neuroticism with 0.71 AUC, and negative affect scales with 0.79 AUC.
DISCUSSION: Our findings demonstrate the feasibility of remotely monitoring cognitive status, social isolation, neuroticism, and psychological well-being. Speech and language patterns were more useful for quantifying cognitive impairment, whereas facial expression and cardiovascular patterns using remote photoplethysmography were more useful for quantifying personality and psychological well-being.
keywords:
Mild Cognitive Impairment, Artificial Intelligence, Psychological Well-being, Social Isolation, Remote Interviews1 INTRODUCTION
@@
2 METHODS
2.1 Participants and Interview Protocol
@@
The study faced the COVID-19 pandemic during the trial recruitment, and cognitive tests were administered by telephone during the pandemic. MoCA was changed to Telephone MoCA during this period. Out of 94 subjects randomized into the intervention group, 52 subjects had in-person MoCA (as opposed to Telephone MoCA). Among them, 39 participants with all the available data, including transcribed data, personality, and NIH-Toolbox emotional battery assessment (discussed later), were used in the current study. The demographic characteristics of these participants are shown in Table 1.
2.2 Outcomes and Clinical Assessment
@@
2.3 Predictors and Multimodal Analysis System for Remote Interview
@@ 3 Results
3.1 Quantifying Cognitive Assessment
@@
3.2 Quantifying Social Network and Psychological Well-being Assessment
@@
4 Discussion
4.1 Identifying Cognitive States.
@@ 4 Discussion
4.2 Quantifying Social Network and Psychological Well-being Assessment.
@@
4.3 Limitations & Future work
@@ Acknowledgments
Hyeokhyen Kwon, Salman Seyedi, Bolaji Omofojoye, and Gari Clifford are partially funded by the National Institute on Deafness and Other Communication Disorders (grant # 1R21DC021029-01A1). Hyeokhyen Kwon and Gari Clifford are also partially supported by the James M. Cox Foundation and Cox Enterprises, Inc., in support of Emory’s Brain Health Center and Georgia Institute of Technology. Gari Clifford is partially supported by the National Center for Advancing Translational Sciences of the National Institutes of Health (NIH) under Award Number UL1TR002378. Gari Clifford and Allen Levey are partially funded by NIH grant #R56AG083845 from the National Institute on Aging. Hirko Dodge is funded by NIH grants and serves as the CEO of the I-CONNECT Foundation, a 501(c)(3) non-profit organization. The I-CONECT study received funding from the NIH: R01AG051628 and R01AG056102. The authors extend their gratitude to the participants of the I-CONECT study.

### v2

@@ Abstract
The aging society urgently requires scalable methods to monitor cognitive decline and identify social and psychological factors indicative of dementia risk in older adults. Our machine learning (ML) models captured facial, acoustic, linguistic, and cardiovascular features from 39 individuals with normal cognition or Mild Cognitive Impairment derived from remote video conversations and classified cognitive status, social isolation, neuroticism, and psychological well-being. Our model could distinguish Clinical Dementia Rating Scale (CDR) of 0.5 (vs. 0) with 0.78 area under the receiver operating characteristic curve (AUC), social isolation with 0.75 AUC, neuroticism with 0.71 AUC, and negative affect scales with 0.79 AUC. Recent advances in machine learning offer new opportunities to remotely detect cognitive impairment and assess associated factors, such as neuroticism and psychological well-being. Our experiment showed that speech and language patterns were more useful for quantifying cognitive impairment, whereas facial expression and cardiovascular patterns using photoplethysmography (PPG) were more useful for quantifying personality and psychological well-being.
Introduction
@@
Materials and methods
Ethics Statement
The I-CONECT study received approval from the Institutional Review Board (IRB) at Oregon Health & Science University (OHSU) (IRB STUDY00015937) through a single IRB process. The patients/participants provided their written informed consent to participate in this study, which also allowed for secondary data analysis. The I-CONECT trial began recruiting participants and collecting data in July 2018. In the second year of recruitment, the COVID-19 pandemic struck the USA. The trial ran for approximately 19 months before the pandemic, from July 2018 to March 2020. The secondary analysis of the dataset from the I-CONECT study was approved by IRB at Emory University (IRB STUDY00007231), as the dataset included identifiable information, such as videos and audio. The IRB was approved on January 8th, 2024, when the study team at Emory University started accessing the dataset for the proposed work.
Participants and Interview Protocol
@@
Cognitive tests were administered by telephone to participants engaging in the study during the COVID-19 pandemic. MoCA was changed to Telephone MoCA during this period. Out of 94 subjects randomized into the intervention group, 52 subjects had in-person MoCA (as opposed to Telephone MoCA). Among them, 39 participants with all the available data, including transcribed data, personality, and NIH-Toolbox emotional battery assessment (discussed later), were used in the current study. The demographic characteristics of these participants are shown in Table 1.
Outcomes and Clinical Assessment
@@
Predictors and Multimodal Analysis System for Remote Interview
@@ Results
Quantifying Cognitive Assessment
@@
Quantifying Social Network and Psychological Well-being Assessment
@@
Discussion
Identifying Cognitive States.
@@ Discussion
Quantifying Social Network and Psychological Well-being Assessment.
@@
Limitations & Future work
@@
Conclusion
With the increasing prevalence of Alzheimer’s disease in an aging society, it is important to quantify the cognitive impairment and psychological well-being of older adults, which helps to facilitate appropriate interventions to slow down cognitive decline [108, 9]. The disparity and the lack of supply in dementia care services significantly challenge the early detection and monitoring of MCI [28, 29]. To tackle this challenge, this work proposed a remote interview analysis that can potentially transform telehealth platforms into behavior monitoring systems for MCI in resource-limited communities. To the best of our knowledge, this is the first work studying the utility of facial, acoustic, linguistic, and cardiovascular patterns independently or combined, which are captured from daily conversations and semi-structured interviews, for quantifying gold-standard ratings of cognitive impairment and psychological well-being, including social satisfaction or negative affect.
Our experiment showed that acoustic and language patterns were most useful for quantifying cognitive impairments and social isolation, indicating the use of telephone-based monitoring for communities with limited internet connectivity. We could also quantify other psychological measures linked to various mental health problems, such as neuroticism or negative affect, through facial patterns and cardiovascular measures from videos. This work makes an important step toward understanding the psychological well-being of older adults through a telehealth platform. Our future work will explore the extent of these technologies for the older adults and MCI population on longitudinal monitoring scenarios and the potential biases in the model due to varying demographic backgrounds [109].
@@ Acknowledgments
Hyeokhyen Kwon, Salman Seyedi, Bolaji Omofojoye, and Gari Clifford are partially funded by the National Institute on Deafness and Other Communication Disorders (grant 1R21DC021029-01A1). Hyeokhyen Kwon and Gari Clifford are also partially supported by the James M. Cox Foundation and Cox Enterprises, Inc., in support of Emory’s Brain Health Center and Georgia Institute of Technology. Gari Clifford is partially supported by the National Center for Advancing Translational Sciences of the National Institutes of Health (NIH) under Award Number UL1TR002378. Gari Clifford and Allen Levey are partially funded by NIH grant R56AG083845 from the National Institute on Aging. Hirko Dodge is funded by NIH grants and serves as the CEO of the I-CONNECT Foundation, a 501(c)(3) non-profit organization. The I-CONECT study received funding from the NIH: R01AG051628 and R01AG056102. The authors extend their gratitude to the participants of the I-CONECT study.